I think of a number. I multiply it by 5. I then subtract 19. Finally, I double the result. The result is 22.
Find the original number.


Kindly, help me with the answer please
THANKS

Found 3 solutions by Boreal, mananth, ikleyn:
Answer by Boreal(15235)

number is x
5x
5x-19
2(5x-19)=22
10x-38=22
x=6

Answer by mananth(16946)

I think of a number. (x)
I multiply it by 5.
5x
I then subtract 19.
5x-19
Finally, I double the result.
(5x-19)*2
The result is 22.
(5x-19)*2=22
5x-19 = 11 (dividing by 2)
5x =19+11
5x =30
x=15
15 the original number.

Answer by ikleyn(52846)

.

Anyone,  who will take a job/(a labor)  to check the solution by @mananth,  will see in a minute
that his answer is  TOTALLY  WRONG.


Use the solution and the answer by  @Boreal,  which is  CORRECT.


Ignore the post by  @mananth,  for your safety . . .